Consider the following information:

MPC = 0.9

Autonomous Investment = R200 m

Autonomous Consumption = R70 m


The equilibrium level of income is:

  1. R300 m
  2. R700 m
  3. R2 000 m
  4. R2 700 m
  5. Impossible to calculate from the information given.
1
Expert's answer
2021-09-02T16:40:17-0400

Marginal propensity to consume: It measures the proportion of the change in the consumption level because of the change in income of the consumer.

It is calculated using the below-mentioned formula;

"MPC =\frac{ ∆C}

{∆Y}"

Here ∆C is the change in the consumption

∆Y is the change  in the income


Given that,

MPC = 0.9

Autonomous Investment = R200 m

Autonomous Consumption = R70 m

At equilibrium level,

National income = Autonomous Investment + Autonomous Consumption 

            "= R200 m + R70 m\\

 = R270 m"

So equilibrium level of income"= \frac{National\space income} 

{MPC}"

So equilibrium level of income "= \frac{270}{

0.9}



 = R300 m"


Based on the above reasoning we can conclude that option 1 "R300 m" is the correct option.